OpenStudy (anonymous):

\[3x \over x-4 minus x +5 over x +4

OpenStudy (anonymous):

\[\frac{3x}{x-4}-\frac{x+5}{x+4}\]

OpenStudy (anonymous):

\[\frac{3x(x+4)-(x-4)(x+5)}{(x-4)(x+4)}\] then algebra in the numerator

OpenStudy (anonymous):

\[\frac{3x(x+4)-(x-4)(x+5)}{(x-4)(x+4)}=\frac{3x^2+12x-(x^2+x-20)}{x^2-16}\] \[\frac{2x^2+11x+20}{x^2-16}\]